a body of mass 0.1 kg drops from a height of 8m onto a hard floor bounces to a height of 2m. Calculate the momentum if the body is in contact with the floor for 0.1 seconds. What is the force exerted on the body (g=ms-²)

1
Expert's answer
2021-11-15T14:00:33-0500

Let's first find the velocity of the body just before it hits a floor. Let's choose the downwards as positive direction. Then, we get:


"v^2=v_0^2+2gh,""v=\sqrt{2gh}=\sqrt{2\times10\ \dfrac{m}{s^2}\times8\ m}=12.65\ \dfrac{m}{s}."

Then, we can find the initial velocity of the body when it bounces from the floor. Let's choose the upwards as the positive direction. Then, we get:


"v^2=v_0^2+2gh,""0=v_0^2+2gh,""v_0=\sqrt{-2gh}=\sqrt{-2\times(-10\ \dfrac{m}{s^2})\times2\ m}=6.32\ \dfrac{m}{s}."

We can find the momentum as follows:


"\Delta p=m(v_f-v_i),""\Delta p=0.1\ kg\times(6.32\ \dfrac{m}{s}-(-12.65\ \dfrac{m}{s}))=1.897\ \dfrac{kg\times m}{s}."

Finally, we can find the force exerted on the body:


"F=\dfrac{\Delta p}{\Delta t}=\dfrac{1.897\ \dfrac{kg\times m}{s}}{0.1\ s}=18.97\ N."